OpenStudy (anonymous):

the surface area of a cylinder is 289cm^2 the height is 21cm what is the diameter

OpenStudy (jack1):

do u know the relationship between radius and surface area?

OpenStudy (anonymous):

no

OpenStudy (jack1):

surface area is the area of the outside: so for a cylinder imagine us split it up the length and unfold it you'll be left with 2 circles and a rectangle

OpenStudy (jack1):



OpenStudy (jack1):

surface area equals sums of area of each shape so area of a circle equation = and area of a rectangle eqn =

OpenStudy (jack1):



OpenStudy (jack1):

reply on a drawing, top right when mouse over pic

OpenStudy (jack1):

circumference of a circle = 2 pi r ... u with me still trevor?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

cool, then: so area of a circle equation = ??? and area of a rectangle eqn = ??? @trevor215

OpenStudy (anonymous):

7.5=area of the circle

OpenStudy (anonymous):

no sorry i got mixed up with another problem thats not the area

OpenStudy (jack1):

so what are your equations to work out the area of a circle?

OpenStudy (jack1):

and area of a rectangle?

OpenStudy (jack1):

k... ill get u started: Equation for area of a rectangle is: Area = Length x Width now u go: circle

OpenStudy (jack1):

Equation for area of a rectangle is: Area = Length x Width in this case, the Length = height of cylinder = 21 and the width = the same as the circumference of one of the circles = 2*pi*r so area of the rectangle = 21 * 2 * pi * r now what are the area's of the circles @trevor215 ?

OpenStudy (anonymous):

idk how to get the radius

OpenStudy (jack1):

ignore the radius for now, just leave it as r so what's the equation for area of a circle

OpenStudy (anonymous):

4 pi r 2